OpenStudy (anonymous):

two forces of 550 newtons and 405 newtons act on a point. the resultant force is 925 newtons find the angle between the two forces

OpenStudy (anonymous):

\[F_1*F_2 *cos(\theta)=F_r\]

OpenStudy (anonymous):

find \[\theta\]